How do i determine the equilibrium level of income using the income expenditure approach and savings investmet approach given the following information;

C=50+7/10, I=50, T=25, G=26, X=17, M=2+1/10Y
1
Expert's answer
2020-06-29T17:15:48-0400

"\text{Using the income expenditure approach}"

"\text{Y=C+I+G+X-M}"

"\text{Y=50+0.7(y-25)+50+26+17-(2+0.1Y)}"

"\text{Y=50+0.7Y-17.5+50+26+17-2+0.1Y}"

"\text{Y=123.5+0.8Y}"

"\text{Y-0.8Y=123.5}"

"\frac{O.2Y}{0.2}" "=\frac{123.5}{0.2}"

"Y=617.5"

"\text{Using savings investment approach}"

"\text{S+T+M}=\text{1+X+C}"

"\text{50+25+{7}/{10}}""Y" "=\text{50+17+26}"

"\text{75+0.1Y}=\text{93-75}"

"\text{0.1Y=18}"

"\text{Y=180}"